OpenStudy (anonymous):

say a function f(num) returns the no. of digits in a number then f(num) + f(num^2) = 1000 is this possible??

OpenStudy (anonymous):

here num is an integer...

OpenStudy (anonymous):

no if you have num =31 then 31+ 961 = 992 if num =32 then 32 + 1024 > 1000

OpenStudy (rsadhvika):

not possible

OpenStudy (rsadhvika):

case I f(num) = k, f(num^2) = 2k case II f(num) = k, f(num^2) = 2k-1

OpenStudy (rsadhvika):

if you solve for k in both cases, you wont get an integer solution so its not possible

OpenStudy (anonymous):

its equal to solving this equation \[\left[ x \right]+\left[ 2x \right]=1000\] where x=log(num) there is no solution for this equation